# ExPerT: Effective and Explainable Evaluation of Personalized Long-Form Text Generation

### v1

@@ 1 Introduction
Figure 1: The ExPerT pipeline.
@@ 3.1 Experimental Setup
We use metrics with publicly available implementations with Python and PyTorch. We use both term-matching and semantic-matching metrics. For term-matching, we employ METEOR Banerjee and Lavie (2005), BLEU Papineni et al. (2002) and ROUGE Lin (2004), which are based on n-gram matching. For semantic-matching, we use BERTScore Zhang et al. (2020), which measures similarity between the representations of the generated and reference text, produced using a text encoder like BERT Devlin et al. (2019) or RoBERTa Liu et al. (2019). [^3] We use the default model recommended for English: https://huggingface.co/FacebookAI/roberta-large. Additionally, we use GEMBA Kocmi and Federmann (2023), which prompts an LLM to generate a score for a given generated output and reference based on predefined criteria. Similarly, G-Eval Liu et al. (2023) performs the same but averages the scores weighted by the probability assigned to each score by the LLM. For both, we employ the prompt shown in Figure 7 in Appendix B using Gemma 2 with 27 billion parameters same as our metric’s LLM. [^4] The model can be found at: https://huggingface.co/google/gemma-2-27b-it The implementations details are explained in Appendix B.
@@ 3.2 Main Findings
We use OpenAI GPT-4o models and Gemma 2 models as the LLMs in ExPerT to investigate this. The results of this experiment are reported in Figure 4. The results show that for smaller LLMs, open-source models (Gemma 2B and 9B) exhibit a smaller alignment with human judgment compared to GPT-4o-mini (0.61 vs 0.64). However, for larger models (Gemma 27B and GPT-4o), both show the same alignment with human judgment (both 0.74). This suggests that for sufficiently large models, there is no significant difference between open-source and proprietary LLMs in terms of alignment with human judgment when used with ExPerT.
@@ 4 Related Work
has been extensively studied for text generation tasks such as machine translation and text summarization Celikyilmaz et al. (2021). Metrics for text evaluation fall into two categories: 1) reference-based and 2) reference-free. Reference-based metrics, such as Exact Match Petroni et al. (2021); Salemi et al. (2023a, b); Salemi and Zamani (2024b, d, c); Kwiatkowski et al. (2019), BLEU Papineni et al. (2002), ROUGE Lin (2004), and METEOR Banerjee and Lavie (2005), rely on n-gram overlap, while more recent approaches like BERTScore Zhang et al. (2020) and BLEURT Sellam et al. (2020) leverage contextual embeddings and learned scoring models. Recent LLM-based methods like GEMBA Kocmi and Federmann (2023), G-Eval Liu et al. (2023), and INSTRUCTSCORE Xu et al. (2023) use LLMs for scoring, often incorporating explanations and predefined criteria for multi-dimensional assessments, such as in UniEval Zhong et al. (2022). Reference-free methods, including LLMs as judges Que et al. (2024); Zheng et al. (2023) and human-LLM collaborations Li et al. (2023b), have also emerged but face challenges like biased evaluation Chen et al. (2024); Stureborg et al. (2024). We utilize LLMs to evaluate personalized text generation with reference outputs, aiming to enhance explainability and alignment with user expectations.
@@
is challenging, as only the user can truly assess whether a response meets their preferences Wang et al. (2023). In automatic evaluation, direct user feedback is not feasible. Previous reference-based methods Salemi et al. (2024b); Kumar et al. (2024); Li et al. (2023a) used n-gram based metrics like ROUGE, BLEU, and METEOR, but these fail to capture nuances like individual preferences, style, or context. Reference-free approaches Wang et al. (2023, 2024) have explored using LLMs to infer user preferences, but they may struggle with accuracy, as they rely on the model’s assumptions, which may not align with the user’s true intentions Dong et al. (2024). This paper aims to improve LLM utilization for evaluating personalized text generation in reference-based scenarios by better capturing content and style similarities to the expected user output and providing explanations about the evaluation process.
@@ Acknowledgment
This work was supported in part by the Center for Intelligent Information Retrieval, in part by the NSF Graduate Research Fellowships Program (GRFP) Award #1938059, in part by Google, and in part by Microsoft. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect those of the sponsor.
@@
This paper focuses exclusively on personalized text generation; however, the proposed approach is generalizable and can be applied to other text generation tasks, such as translation and summarization. Exploring these broader applications is beyond the scope of this work and is left for the future research.
@@ Appendix A Datasets & Task Definition
This paper utilizes the LongLaMP benchmark Kumar et al. (2024) for our experiments. Each example in each dataset corresponds to a unique user and includes: (1) an input prompt relevant to the task, (2) an expected output personalized for the user, and (3) a user profile containing historical data, such as previously generated texts, to reflect the user’s writing style and preferences. Our experiments are conducted using the user-based setting of the LongLaMP benchmark. The dataset statistics are provided in Table 2. The benchmark includes three [^6] The LongLaMP benchmark originally consists of four personalized generation tasks. However, due to privacy concerns regarding the email dataset and licensing issues for human annotation, we exclude that task. personalized long-form generation tasks:
@@ Appendix B Baselines Details
Papineni et al. (2002) is a widely used metric for evaluating the quality of machine-generated text. It measures the overlap between n-grams of the generated text and one or more reference texts, focusing on precision to determine how much of the generated output matches the references. BLEU employs a brevity penalty to discourage excessively short translations and calculates a geometric mean of precision scores across different n-gram sizes. We utilize the HuggingFace implementation of this metric. [^7] This metric can be found at: https://huggingface.co/spaces/evaluate-metric/bleu
@@
Lin (2004) is a metric designed to evaluate text generation tasks by comparing the overlap of the longest common subsequences between a generated text and reference. ROUGE-L emphasizes on sequential relationship of words, capturing structural similarity. We utilize the HuggingFace implementation of this metric. [^8] This metric can be found at: https://huggingface.co/spaces/evaluate-metric/rouge
@@
Banerjee and Lavie (2005) is a widely used automatic evaluation metric designed to assess the quality of a generated output by comparing them to a reference. Instead of relying primarily on exact n-gram matches, METEOR incorporates stemming, synonym matching, and a flexible alignment approach to capture variations in word usage and sentence structure. We utilize the HuggingFace implementation of this metric. [^9] This metric can be found at: https://huggingface.co/spaces/evaluate-metric/meteor
@@
Zhang et al. (2020) is a metric for evaluating text generation tasks that leverages contextualized embeddings from pre-trained models like BERT. Unlike traditional n-gram-based metrics, BERTScore computes similarity based on the cosine similarity of word embeddings, capturing semantic meaning rather than exact word matches. It uses token-level embeddings to compare each word in the generated text with its corresponding word in the reference, considering both precision and recall. This allows BERTScore to assess the quality of generated texts more effectively, especially when dealing with synonyms. We utilize the HuggingFace implementation of this metric. [^10] This metric can be found at : https://huggingface.co/spaces/evaluate-metric/bertscore
@@
Kocmi and Federmann (2023) is a metric for evaluating text generation tasks that utilizes LLMs with predefined evaluation criteria. It compares the generated text in response to a prompt with a reference output for the same prompt to assess the quality of the generated text. In this approach, the prompt, generated text, expected output, and a predefined evaluation criterion are provided to an LLM. The model is then asked to generate a score for the generated output by comparing it to the reference, taking the specified criteria into account. In this paper, we utilize the pointwise scoring prompt shown in Figure 7 to generate the scores. We set the model’s temperature to zero to obtain more deterministic results. Additionally, we limit the consideration to a maximum of 512 tokens from both the generated output and the expected output. For backbone LLM, we utilize an instruction-tuned Gemma 2 Gemma-Team (2024) with 27 billion parameters [^11] The model can be found at: https://huggingface.co/google/gemma-2-27b-it using the VLLM library Kwon et al. (2023). [^12] This framework can be found at: https://github.com/vllm-project/vllm
@@
Liu et al. (2023) is another LLM-based metric for text generation evaluation, similar to GEMBA, which takes an input prompt, generated output, and reference output along with predefined criteria to score the generated output. However, G-Eval considers the probability of each score in the final score calculation. Specifically, the model multiplies each score in the predefined criteria by the probability that the model assigns to that score, then calculates a weighted average of the scores as the final score. To implement this, following the original paper, we generate 20 scores using the LLM with a high temperature of 1. Based on the count of each score, we calculate the probabilities for each score. We then average the scores based on these probabilities to obtain the final score. In this paper, we utilize the pointwise scoring prompt shown in Figure 7 to generate the scores. Additionally, we limit the consideration to a maximum of 512 tokens from both the generated output and the expected output. For the backbone LLM, we use an instruction-tuned Gemma 2 Gemma-Team (2024) with 27 billion parameters [^13] The model can be found at: https://huggingface.co/google/gemma-2-27b-it with the VLLM library Kwon et al. (2023). [^14] This framework can be found at: https://github.com/vllm-project/vllm
@@ Appendix C Personalizing LLMs through RAG
In this paper, we personalize both GPT-4o-mini [^15] This model is not open source and is served by OpenAI and described at: https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/ and Gemma 1.1161616The checkpoint can be found at: google/gemma-1.1-2b-it. For GPT-4o-mini, we apply the method described earlier to generate personalized responses. In contrast, for Gemma 1.1, we fine-tune the model on the LongLaMP benchmark to adapt it to personalized text generation tasks. For fine-tuning, we use a sequence-to-sequence loss function Sutskever et al. (2014). Given the personalized prompt $x_{u}$ produced using the method described earlier, the model is trained to generate the expected output $y$. This ensures that the model learns to generate personalized responses based on the input prompt tailored to the user’s profile. We train the model for 5000 steps using a multi-tasking approach across all datasets in the LongLaMP benchmark. The training is conducted with a learning rate of $5\times 10^{-5}$, using the Adam optimizer Kingma and Ba (2015) with a weight decay of $10^{-4}$, and a batch size of 64. We perform 250 warmup steps to stabilize training. The model’s context length is set to 2048 tokens, and we limit each retrieved document to the first 400 tokens when generating the personalized prompt. For inference, we set the temperature to 0.1 to ensure more deterministic output generation.

### v2

@@ 1 Introduction
Figure 1: ExPerT pipeline: Generated and reference outputs are first decomposed into atomic aspects along with their corresponding evidences. Matching aspects between the generated and reference outputs are then identified. Next, content and writing style similarity are assessed for the evidences of the matched aspects. Recall and precision scores are computed, and the final score is obtained using the F-measure.
@@ 3.1 Experimental Setup
We use metrics with publicly available implementations with Python and PyTorch. We use both term-matching and semantic-matching metrics. For term-matching, we employ METEOR Banerjee and Lavie (2005), BLEU Papineni et al. (2002) and ROUGE Lin (2004), which are based on n-gram matching. For semantic-matching, we use BERTScore Zhang et al. (2020), which measures similarity between the representations of the generated and reference text, produced using a text encoder like BERT Devlin et al. (2019) or RoBERTa Liu et al. (2019). [^3] We use the default model recommended for English: https://hf.co/FacebookAI/roberta-large. Additionally, we use GEMBA Kocmi and Federmann (2023), which prompts an LLM to generate a score for a given generated output and reference based on predefined criteria. Similarly, G-Eval Liu et al. (2023) performs the same but averages the scores weighted by the probability assigned to each score by the LLM. For both, we employ the prompt shown in Figure 7 in Appendix B using Gemma 2 with 27 billion parameters same as our metric’s LLM. [^4] The model can be found at: https://hf.co/google/gemma-2-27b-it The implementations details are explained in Appendix B.
@@ 3.2 Main Findings
We use OpenAI GPT-4o and Gemma 2 models as the LLMs in ExPerT to investigate this. The results of this experiment are reported in Figure 4. The results show that for smaller LLMs, open-source models (Gemma 2B and 9B) exhibit a smaller alignment with human judgment compared to GPT-4o-mini (0.61 vs 0.64). However, for larger models (Gemma 27B and GPT-4o), both show the same alignment with human judgment (both 0.74). This suggests that for sufficiently large models, there is no significant difference between open-source and proprietary LLMs in terms of alignment with human judgment when used with ExPerT.
@@
How efficient is ExPerT compared to the LLM-based baselines?
To enable a standardized cost comparison, we define a single invocation of the language model (LLM) as one unit of cost. Under this definition, any metric that queries the LLM once incurs a cost of 1. GEMBA, by design, has a fixed cost of 1, whereas the cost of G-Eval corresponds to the total number of LLM calls needed to compute individual component scores and aggregate them via a weighted average. In our experiments, G-Eval required 20 LLM calls per instance. For ExPerT, the number of LLM invocations varies based on the number of aspects and concepts identified in both the expected and generated outputs. In our evaluation on 100 samples from the human-annotated dataset, we observed that ExPerT makes an average of 18.6 LLM calls per instance. These calls are used for extracting aspects from both outputs and aligning them in terms of content and style. This analysis suggests that ExPerT provides a more cost-efficient and robust evaluation compared to G-Eval. While GEMBA is minimal in cost, requiring only one LLM call, it is substantially more susceptible to adversarial inputs and demonstrates reduced evaluation reliability.Overall, ExPerT provides a balanced trade-off among effectiveness, robustness, reliability, and computational efficiency when compared to existing LLM-based evaluation methods.
@@ 4 Related Work
has been extensively studied for tasks such as machine translation and summarization Celikyilmaz et al. (2021). Metrics for text evaluation fall into two categories: 1) reference-based and 2) reference-free. Reference-based metrics, such as Exact Match Petroni et al. (2021); Salemi et al. (2023a, b); Salemi and Zamani (2024b, d, c); Kwiatkowski et al. (2019), BLEU Papineni et al. (2002), ROUGE Lin (2004), and METEOR Banerjee and Lavie (2005), rely on n-gram overlap, while more recent approaches like BERTScore Zhang et al. (2020) and BLEURT Sellam et al. (2020) leverage contextual embeddings and learned scoring models. Recent LLM-based methods like GEMBA Kocmi and Federmann (2023), G-Eval Liu et al. (2023), and INSTRUCTSCORE Xu et al. (2023) use LLMs for scoring, often incorporating explanations and predefined criteria for multi-dimensional assessments, such as in UniEval Zhong et al. (2022). Reference-free methods, including LLMs as judges Que et al. (2024); Zheng et al. (2023) and human-LLM collaborations Li et al. (2023b), have also emerged but face challenges like biased evaluation Chen et al. (2024); Stureborg et al. (2024). We utilize LLMs to evaluate personalized text generation with reference outputs, aiming to enhance explainability and alignment with user expectations.
@@
is challenging, as only the user can truly assess whether a response meets their preferences Wang et al. (2023). In automatic evaluation, direct user feedback is not feasible. Previous reference-based methods Salemi et al. (2024b); Kumar et al. (2024); Li et al. (2023a) used n-gram based metrics like ROUGE, BLEU, and METEOR, but these fail to capture nuances like individual preferences, style, or context. Furthermore, the use of rubric-based methods with a personalized-trained network has been explored Hashemi et al. (2024). However, this approach relies on user-specific training data for each questions in the rubric, which is not readily available in many real-world scenarios and cannot be a baseline in our experiments. Reference-free approaches Wang et al. (2023, 2024) have explored using LLMs to infer user preferences, but they may struggle with accuracy, as they rely on the model’s assumptions, which may not align with the user’s true intentions Dong et al. (2024). This paper aims to improve LLM utilization for evaluating personalized text generation in reference-based scenarios by better capturing content and style similarities to the expected user output and providing explanations about the evaluation process.
@@ Acknowledgment
This work was supported in part by the Center for Intelligent Information Retrieval, in part by NSF grant #2143434, in part by the NSF Graduate Research Fellowships Program (GRFP) Award #1938059, in part by Google, and in part by Microsoft. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect those of the sponsor.
@@
This paper focuses exclusively on personalized text generation; however, the proposed approach is generalizable and can be applied to other text generation tasks, such as machine translation and summarization. Investigating these broader applications is beyond the scope of this work and is left for future research. Additionally, to the best of our knowledge, the LongLaMP benchmark is the only benchmark for long-form personalized text generation in a reference-based setting. Evaluating the effectiveness of this metric in other personalized text generation tasks not covered by this benchmark could provide valuable information.
Extension to Other Languages.
This paper focuses exclusively on the English language, as, to the best of our knowledge, no datasets are available for studying personalization in other languages. Nonetheless, extending this research to other languages could yield valuable insights.
@@ Appendix A Datasets & Task Definition
This paper utilizes the LongLaMP benchmark Kumar et al. (2024) for our experiments, which is a publicly accessible dataset for long-form personalized text generation. [^6] This benchmark does not specify any licensing restrictions, so we utilized it solely for research purposes in accordance with its intended use. Each example in each dataset corresponds to a unique user and includes: (1) an input prompt relevant to the task, (2) an expected output personalized for the user, and (3) a user profile containing historical data, such as previously generated texts, to reflect the user’s writing style and preferences. Our experiments are conducted using the user-based setting of the LongLaMP benchmark. The dataset statistics are provided in Table 2. The benchmark includes three [^7] The LongLaMP benchmark originally consists of four personalized generation tasks. However, due to privacy concerns regarding the email dataset and licensing issues for human annotation, we exclude that task. personalized long-form generation tasks:
@@ Appendix B Baselines Details
Papineni et al. (2002) is a widely used metric for evaluating the quality of machine-generated text. It measures the overlap between n-grams of the generated text and one or more reference texts, focusing on precision to determine how much of the generated output matches the references. BLEU employs a brevity penalty to discourage excessively short translations and calculates a geometric mean of precision scores across different n-gram sizes. We utilize the HuggingFace implementation of this metric. [^8] This metric can be found at: https://hf.co/spaces/evaluate-metric/bleu
@@
Lin (2004) is a metric designed to evaluate text generation tasks by comparing the overlap of the longest common subsequences between a generated text and reference. ROUGE-L emphasizes on sequential relationship of words, capturing structural similarity. We utilize the HuggingFace implementation of this metric. [^9] This metric can be found at: https://hf.co/spaces/evaluate-metric/rouge
@@
Banerjee and Lavie (2005) is a widely used automatic evaluation metric designed to assess the quality of a generated output by comparing them to a reference. Instead of relying primarily on exact n-gram matches, METEOR incorporates stemming, synonym matching, and a flexible alignment approach to capture variations in word usage and sentence structure. We utilize the HuggingFace implementation of this metric. [^10] This metric can be found at: https://hf.co/spaces/evaluate-metric/meteor
@@
Zhang et al. (2020) is a metric for evaluating text generation tasks that leverages contextualized embeddings from pre-trained models like BERT. Unlike traditional n-gram-based metrics, BERTScore computes similarity based on the cosine similarity of word embeddings, capturing semantic meaning rather than exact word matches. It uses token-level embeddings to compare each word in the generated text with its corresponding word in the reference, considering both precision and recall. This allows BERTScore to assess the quality of generated texts more effectively, especially when dealing with synonyms. We utilize the HuggingFace implementation of this metric. [^11] This metric can be found at : https://hf.co/spaces/evaluate-metric/bertscore
@@
Kocmi and Federmann (2023) is a metric for evaluating text generation tasks that utilizes LLMs with predefined evaluation criteria. It compares the generated text in response to a prompt with a reference output for the same prompt to assess the quality of the generated text. In this approach, the prompt, generated text, expected output, and a predefined evaluation criterion are provided to an LLM. The model is then asked to generate a score for the generated output by comparing it to the reference, taking the specified criteria into account. In this paper, we utilize the pointwise scoring prompt shown in Figure 7 to generate the scores. We set the model’s temperature to zero to obtain more deterministic results. Additionally, we limit the consideration to a maximum of 512 tokens from both the generated output and the expected output. For backbone LLM, we utilize an instruction-tuned Gemma 2 Gemma-Team (2024) with 27 billion parameters [^12] The model can be found at: https://hf.co/google/gemma-2-27b-it using the VLLM library Kwon et al. (2023). [^13] This framework can be found at: https://github.com/vllm-project/vllm
@@
Liu et al. (2023) is another LLM-based metric for text generation evaluation, similar to GEMBA, which takes an input prompt, generated output, and reference output along with predefined criteria to score the generated output. However, G-Eval considers the probability of each score in the final score calculation. Specifically, the model multiplies each score in the predefined criteria by the probability that the model assigns to that score, then calculates a weighted average of the scores as the final score. To implement this, following the original paper, we generate 20 scores using the LLM with a high temperature of 1. Based on the count of each score, we calculate the probabilities for each score. We then average the scores based on these probabilities to obtain the final score. In this paper, we utilize the pointwise scoring prompt shown in Figure 7 to generate the scores. Additionally, we limit the consideration to a maximum of 512 tokens from both the generated output and the expected output. For the backbone LLM, we use an instruction-tuned Gemma 2 Gemma-Team (2024) with 27 billion parameters [^14] The model can be found at: https://hf.co/google/gemma-2-27b-it with the VLLM library Kwon et al. (2023). [^15] This framework can be found at: https://github.com/vllm-project/vllm
@@ Appendix C Personalizing LLMs through RAG
In this paper, we personalize both GPT-4o-mini [^16] This model is not open source and is served by OpenAI and described at: https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/ and Gemma 1.1171717The checkpoint can be found at: google/gemma-1.1-2b-it. For GPT-4o-mini, we apply the method described earlier to generate personalized responses. In contrast, for Gemma 1.1, we fine-tune the model on the LongLaMP benchmark to adapt it to personalized text generation tasks. For fine-tuning, we use a sequence-to-sequence loss function Sutskever et al. (2014). Given the personalized prompt $x_{u}$ produced using the method described earlier, the model is trained to generate the expected output $y$. This ensures that the model learns to generate personalized responses based on the input prompt tailored to the user’s profile. We train the model for 5000 steps using a multi-tasking approach across all datasets in the LongLaMP benchmark. The training is conducted with a learning rate of $5\times 10^{-5}$, using the Adam optimizer Kingma and Ba (2015) with a weight decay of $10^{-4}$, and a batch size of 64. We perform 250 warmup steps to stabilize training. The model’s context length is set to 2048 tokens, and we limit each retrieved document to the first 400 tokens when generating the personalized prompt. For inference, we set the temperature to 0.1 to ensure more deterministic output generation.